A compound contains 32% carbon, 4% hydrogen and rest oxygen.
Its vapour density is 75. Calculate the empirical and molecular formula.
Given Answer :  C2H3O3, C4H6O6
the procedure of solving is not given and the answer is wrong! 

Grade:11

1 Answers

ram charan
7 years ago
calculate the no of moles of each element in compound 
no of moles of carbon=32/12=2.66
no of moles of hydrogen=4/1=4
no of moles of oxygen = 64/16=4
then find ratio of no of moles carbon:hydrogen:oxygen
2.6:4:4=1.3:2:2=1.95:3:3
therefore emperical formula is C2H3O3.